Cairns owns 70 percent of the voting stock of Hamilton, Inc. The parent’s interest was acquired several years ago on the date that the subsidiary was formed. Consequently, no goodwill or other allocation was recorded in connection with the acquisition. Cairns uses the equity method in its internal records to account for its investment in Hamilton. On January 1, 2014, Hamilton sold $1,200,000 in 10-year bonds to the public at 110. The bonds had a cash interest rate of 9 percent payable every December 31. Cairns acquired 40 percent of these bonds at 92 percent of face value on January 1, 2016. Both companies utilize the straight-line method of amortization. Prepare the consolidation worksheet entries to recognize the effects of the intra-entity bonds at each of the following dates. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.) December 31, 2016 December 31, 2017 December 31, 2018
Answer #1

Amortisation value of bond by Hamilton : 1,200,000/10 = 120,000.

Face Value of Bond as on 1.1.2016 : 1200,000- (12000*2= 960000

40% acquired by Cairne ie 40%*960000 = 384000. Price paid 384000*92% = 353,280.

Amortisation by Cairne: 353280/8 = 44160 per year.

40% of Security Premium: 120,000*40% = 48000.

  • 31.12.16 :
    • 1) Interest Received Dr 34560 (960,000*9%*40%)
    • To Interest Paid 34560
    • 2) 9% Bonds (960,000*40%) Dr 384000
    • Security premium Dr 48000
    • TO Investment in Hamilton Bonds 353280
    • To Capital Reserve 78,720
    • .
    • 3) Ammortisation of Bonds Dr 48000
    • To Amortisation of Investment 44160
    • To General Reserve 3840
  • 31.12.17
    • 1) Internet Received Dr 30240 (840000*9%*40%)
    • TO Interest Paid 30240
    • .
    • 2) (Same as entry 3 on 31.12.16)
  • 31.12.18
    • 1) Interest Received Dr 25,920 ( 720,00*9%*40%)
    • To interest Paid 25920
    • .
    • 2) Same as entry 3 on 31.12.16